# Experimental Design for Active Transductive Inference in Large Language Models

### v1

@@ Appendix D Prompt Examples
Iris Classification Dataset: Below we give an example of how we use the prompts to be used in the LLM for the Iris misclassification task. Similar types of prompt can be found in Dinh et al. (2022); Suzgun et al. (2022). In the following prompt, the point $\mathbf{x}_{i}$ chosen by GO at trial $1$, time $1$ is given by $[42,29,69,75]\in\mathbb{R}^{4}$. The label of this point is done by human and is the class of this data point (shown as $2$). Finally, this prompt is fed into the LLM as $x1=42,x2=29,x3=69,x4=75$ for the target query $\mathbf{x}_{*,1}$ represented as $x1=69,x2=33,x3=64,x4=54$ for the LLM prediction of the class of the target query. Finally observe that we proceed in the similar way for each of the $\mathbf{x}_{*,2}$ as $x1=19,x2=58,x3=8,x4=4$, $\mathbf{x}_{*,2}$ as $x1=19,x2=62,x3=5,x4=8$, and $\mathbf{x}_{*,4}$ as $x1=50,x2=25,x3=78,x4=54$. Then we move to the next iteration of the trail, and this interaction proceeds till the end of the horizon.
trial 1
action in time 1: [42 29 69 75]
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 69, x2 = 33, x3 = 64, x4 = 54 label is
answer in time 1: 0
True label: 1
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 19, x2 = 58, x3 = 8, x4 = 4 label is
answer in time 1: 1
True label: 0
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 19, x2 = 62, x3 = 5, x4 = 8 label is
answer in time 1: 0
True label: 0
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 42, x2 = 33, x3 = 69, x4 = 96 label is
answer in time 1: 1
True label: 2
final_prompt at time 1:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 69, x2 = 33, x3 = 64, x4 = 54 label is
answer in time 1: 1
True label: 1
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 50, x2 = 25, x3 = 78, x4 = 54 label is
answer in time 2: 1
True label: 2
action in time 2: [56 12 58 50]
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 69, x2 = 33, x3 = 64, x4 = 54 label is
answer in time 2: 0
True label: 1
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 19, x2 = 58, x3 = 8, x4 = 4 label is
answer in time 2: 3
True label: 0
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 19, x2 = 62, x3 = 5, x4 = 8 label is
answer in time 2: 0
True label: 0
final_prompt at time 2:
x1 = 42, x2 = 29, x3 = 69, x4 = 75 label is 2. x1 = 56, x2 = 12, x3 = 58, x4 = 50 label is 1. x1 = 42, x2 = 33, x3 = 69, x4 = 96 label is
answer in time 2: 1
True label: 2
@@
trial 1
action in time 1: [’Frozen (2013)’ ’Brave (2012)’]
final_prompt at time 1:
[’Frozen (2013)’ ’Brave (2012)’] label is 2, Monsters Inc, Frozen (2013):
answer in time 1: 2
True label: 2
action in time 2: [’Wall’ ’Monsters Inc’]
final_prompt at time 2:
[’Wall’ ’Monsters Inc’] label is 2, Monsters Inc, Frozen (2013):
answer in time 2: 2
True label: 2
[…]
trial 1
action in time 1: [’Monty Python and the Holy Grail (1975)’]
final_prompt at time 1:
Monty Python and the Holy Grail (1975) label is 4, Moneyball (2011):
answer in time 0: 4
 True label: 1
final_prompt at time 1:
Monty Python and the Holy Grail (1975) label is 4, Slap Shot (1977):
answer in time 1: 4
 True label: 1
Appendix E Ablation Studies
E.1 Extensive ablation experiments with Greedy-NN
In this section we conduct an ablation study for various versions of Greedy-NN.
Greedy-NN-max-max: This is the version implemented in the paper in all experiments which we call as Greedy-NN. For clarity in this section, we will call this as Greedy-NN-max-max. Recall the Greedy-NN-max-max implemented in Section 5.1 is as follows: The chosen point in round $t$ aligns the most with the target point $\mathbf{x}_{*}$ such that $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}\mathbf{x}_{i}$. For multiple target points $\mathbf{x}_{*,k}$ Greedy-NN-max-max chooses $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\max_{k}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$.
Greedy-NN-max-min: The Greedy-NN-max-min works as follows: The chosen point in round $t$ aligns the most with the target point $\mathbf{x}_{*}$ such that $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}\mathbf{x}_{i}$. For multiple target points $\mathbf{x}_{*,k}$ Greedy-NN-max-min chooses $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\min_{k}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$.
Greedy-NN-max-mean: Before we state this method, we define the notation $\mathrm{avg}(\mathbf{x}_{i})=\frac{1}{K}\sum_{k=1}^{K}\mathbf{x}_{*,k}^{\top}\mathbf{x}_{i}$. Then define the score for a point $\mathbf{x}_{i}$ with respect to $\mathbf{x}_{*}$ and the distance from the average $\mathrm{avg}(\mathbf{x}_{i})$ as $s(\mathbf{x}_{i},\mathbf{x}_{*})=|\mathbf{x}_{*}^{\top}\mathbf{x}_{i}-\mathrm{avg}(\mathbf{x}_{i})|$. Now, the Greedy-NN-max-mean works as follows: At round $t$ choose a point that aligns most with the target point $\mathbf{x}_{*}$. So the point is chosen such that $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathbf{x}_{*}^{\top}\mathbf{x}_{i}$. For multiple target points $\mathbf{x}_{*,k}$ Greedy-NN-max-mean chooses $I_{t}\leftarrow\mathop{\rm arg\,max}_{i\in\mathcal{U}_{t}}\mathop{\rm arg\,min}_{k}s(\mathbf{x}_{i},\mathbf{x}_{*,k})$. Observe that the inner $\min$ chooses the point $\mathbf{x}_{i}$ closest to the average of all the target points $\mathbf{x}_{*,k}$ to balance exploration.
(a) Classification on ARC Expansion and Contraction Experiment
(b) Classification on ARC Rotation Experiment
(c) Classification on PCFG Add Subtract
(d) Classification on Iris
Figure 7: Experiment results with Falcon 40B on ARC, PCFG, classification tasks. For the classification tasks, we show the cumulative mis-classification error in the y-axis.
The experiments with all these variants are shown in Figure 7. From this Figure 7 we see that none of the variants of the Greedy-NN are reducing the misclassification error or 0-1 error as fast as GO or SAL. This performance of the many variants of Greedy-NN can be attributed to its lack of correctly balancing exploration-exploitation.
E.2 Instructor embedding vs numerical features
Instructor embedding can be derived for any inputs that can be represented as strings. In this section, we examine whether Instructor embedding is helpful in active experiment design when numerical input features are considered. As shown in Figure 8, using Instructor embedding for the numerical features always hurts the performance of G-optimal active experiment design.
Figure 8: Comparison of the cumulative error for active experiment design conducted with original numerical features vs. with Instructor embedding. The selected datasets all use numerical input features rather than natural language. GO is used in all the experiments in this figure.
Appendix F Table of Notations
@@
$n$ Total unlabeled points
@@
$\mathbf{x}_{i}$ Feature of sample point $i$
$\mathbf{x}_{*,k}$ $k$-th target point
@@
$Y_{*}$ Target label
$H_{t}=(X_{\ell},Y_{\ell})_{\ell\in[t-1]}$ History of $t-1$ previously labeled points
$p(\cdot\mid\mathbf{x},H_{t})$ Distribution of the label of point $\mathbf{x}$ conditioned on $H_{t}$
@@
$\mathcal{L}_{t}$ Set of labeled points
$\mathcal{U}_{t}$ Set of unlabeled points

### v2

@@ Appendix D Prompt Examples
Classification Dataset Prompts : Below we give an example of how we use the prompts to be used in the LLM for the Iris misclassification task. Similar types of prompts can be found in Dinh et al. [2022], Suzgun et al. [2022]. This is shown in Figure 3(a). Note that since we have the feature representation of the training and test examples from the dataset, we directly use them as $\mathbf{x}_{i}$ and $\mathbf{x}_{*,k}$.
Regression Dataset Prompts : In Figure 3(b) we give an example of a prompt for regression task in Fifa dataset. Note that since we have the feature representation of the training and test examples from the dataset, we directly use them as $\mathbf{x}_{i}$ and $\mathbf{x}_{*,k}$.
(a) Classification Prompt
(b) Regression Prompt
(c) Movie Prompt
(d) Entity Prompt
Figure 3: Prompt examples for Classification, Regression, Movie, and Prompt
(a) Theme Prompt
(b) PCFG Prompt
Figure 4: Prompt examples for Theme and PCFG tasks
(a) ARC Prompt
Figure 5: Prompt examples for ARC task
@@
Movie Name Experiment: We use a similar technique as in Iris dataset for this setting. The labels of the movie genres belong to $5$ classes as follows: romance, horror, thriller, sport, and action. At every iteration we pass a set of test movie name examples where each $\mathbf{x}_{*,k}$ is now movie name. Note that we feed the natural language text to the LLM as prompts as shown in Figure 3(c). However, to run GO, SAL, and other baselines we require a featurization of these natural language prompts. We obtain a $768$ dimensional featurized representation of the movie names using Instructor embedding [Su et al., 2022]. This constitutes $\mathbf{x}_{i}\in\mathbb{R}^{768}$ and $\mathbf{x}_{*,k}\in\mathbb{R}^{768}$.
Entity Name Experiment: The labels of the entity genres belong to $5$ classes as follows: mountains, seas, rivers, vehicles, and celebrities. At every iteration, we pass a set of test entity name examples where each $\mathbf{x}_{*,k}$ is now an entity name. Note that we feed the natural language text to the LLM as prompts as shown in Figure 3(d). However, to run GO, SAL, and other baselines we require a featurization of these natural language prompts. Again, we obtain a $768$ dimensional featurized representation of the entity names using Instructor embedding [Su et al., 2022]. This constitutes $\mathbf{x}_{i}\in\mathbb{R}^{768}$ and $\mathbf{x}_{*,k}\in\mathbb{R}^{768}$.
PCFG Experiment: We show an example of this prompt in Figure 4(b). Here we concatenate the sequence to obtain training examples $\mathbf{x}_{i}$ and test examples $\mathbf{x}_{*,k}$. So a sequence of $4$ integers of length $4$ will be represented by $\mathbf{x}_{i},\mathbf{x}_{*,k}\in\mathbb{R}^{16}$. Similarly the label $Y_{i}$ and $Y_{*,k}$ consist of sequence of $5$ integers of length $4$ which we concatentate to get a vector of length $\mathbb{R}^{20}$.
ARC Experiment: We show an example of this prompt in Figure 5(a). Here we vectorized the $4\times 4$ matrix to obtain training examples $\mathbf{x}_{i}\in\mathbb{R}^{16}$ and test examples $\mathbf{x}_{*,k}\in\mathbb{R}^{16}$. Similarly the label $Y_{i}$ and $Y_{*,k}$ consist of vectorized matrices of length $\mathbb{R}^{16}$.
Appendix E Table of Notations
@@
$n$ Total unlabeled examples
@@
$\mathbf{x}_{i}$ Feature of sample example $i$
$\mathbf{x}_{*,k}$ $k$-th test example
@@
$Y_{*}$ Label
$H_{t}=(X_{\ell},Y_{\ell})_{\ell\in[t-1]}$ History of $t-1$ previously labeled examples
$p(\cdot\mid\mathbf{x},H_{t})$ Distribution of the label of example $\mathbf{x}$ conditioned on $H_{t}$
@@
$\mathcal{L}_{t}$ Set of labeled examples
$\mathcal{U}_{t}$ Set of unlabeled examples